OpenStudy (amtran_bus):

Find a function whose graph is a parabola with vertex (1, −5) and that passes through the point (3, 15).

OpenStudy (dan815):

This parabola is facing up you can tell by the fact that the vextex is below a point

OpenStudy (amtran_bus):

I know how to do it I think.

OpenStudy (dan815):

oh ok

OpenStudy (amtran_bus):

Gotta find the A

ganeshie8 (ganeshie8):

find it

OpenStudy (dan815):

y=ax^2+bx+c y'=2ax+b 0=2ax+b x=-b/2a now u know that x=1, as that is the x value of the vertex so 1=-b/2a -b=2a now you also know that -5 is the value of y there, so 5=a(-b/2a)^2+b(-b/2a)+c <--- sub in -b=2a in there and finally solve for c with the final initial value

OpenStudy (jdoe0001):

\(\bf y={\color{brown}{ a}}(x-h)^2+k\qquad vertex\ ({\color{blue}{ 3,15}})\to y={\color{brown}{ a}}(x-{\color{blue}{ 3}})^2{\color{blue}{ +5}} \\ \quad \\ \textit{what's }"a"?\qquad \textit{passes through (3,15)}\qquad thus \\ \quad \\ {\color{olive}{ 15}}={\color{brown}{ a}}({\color{olive}{ 3}}-{\color{blue}{ 3}})^2{\color{blue}{ +5}}\)